• It costs an airline $100,000 to fly a 200-seat plane across the country. Average cost per seat therefore is $500 ($100,000/200)
  • Suppose it charges each passenger $600 for a plane ticket but on average can only sell 190 tickets. It can sell all 200 tickets if it offers a 50% discount on 10 tickets on a first-come basis to standby passengers. The additional cost of taking in the 10 additional passengers is $20 per passenger for the meals, snack and refreshments.


1.What is the marginal benefit of taking in an additional passenger with the discount?

2.What is the additional cost per additional passenger?

3.What is net profit of the airline if it does not offer the discount?

4.What is the net profit if it offers the discount?

5.Should it offer the discount, yes or no?


1
Expert's answer
2021-04-19T18:48:39-0400

1.The marginal benefit of taking in an additional passenger with the discount is 600/2 = $300.

2.The additional cost per additional passenger is $20.

3.The net profit of the airline if it does not offer the discount is:

TP = 600×190 - (100,000 + 20×190) = 10,200.

4. The net profit if it offers the discount is:

TP = (600×190 + 300×10) - (100,000 + 20×200) = 13,000.

5. It should offer the discount.